Find the perimeter of a rectangular board whose length is 3 feet longer than the width. p=2l+2w
Found 2 solutions by solver91311, ewatrrr:
Answer by solver91311(24713)



If the length is 3 feet longer than the width, then

Then substituting in the formula for perimeter we get:



which simplifies to:




John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by ewatrrr(24785)

Hi
p = 2*l + 2*w (p being the perimeter, l being the length, w the width)
.
the question states the following to be true:
l = w + 3
.
P = 2*(w + 3) + 2*w
.
P = 4*w + 6